# Intelligent Polarforming Antenna Enhanced Sensing and Communication: Modeling and Optimization

### v1

@@ VI Simulation Results
In the simulation, we set $N=4$, $B=16$, $K=30$, and $\mathcal{C}$ as a cube with $A=1$ meter (m) sides. Users are situated within a 3D coverage area, which forms a spherical annulus ranging from $20$ m to $200$ m in radial distance from the BS center. The carrier frequency is 24 GHz. The inter-antenna spacing on each IPA subarray is $\frac{\lambda}{2}$ and the minimum inter-subarray distance is $d_{\min}=\frac{\sqrt{2}}{2}\lambda+\frac{\lambda}{2}$. In Algorithm 3, we set $S=200$, $\kappa^{(i)}={i^{-0.2}}$, and $I_{\mathrm{iter}}=100$. We set the number of channel samples $\overline{L}=4000$ with $L_{\mathrm{S}}=40$, unless specified otherwise.
To start with, we evaluate the performance of the proposed polarforming-based localization algorithm and compare it with the direct positioning method in [34]. The direct positioning method sets $L=T_{\mathrm{s}}$ with $P=1$ and applies maximum likelihood (ML) estimation to iteratively optimize the users’ DoA vectors $\mathbf{f}_{k}$, user-to-BS distances $d_{k}$, and dynamic polarformed channels $\eta_{m,k}$ to maximize the log-likelihood function. The localization error is calculated by $\sqrt{\mathbb{E}\left\{\|\hat{\mathbf{p}}-\mathbf{p}\|^{2}\right\}}$ [35], where $\mathbf{p}$ represents the true locations of all users, and $\hat{\mathbf{p}}$ denotes their estimated locations. During the localization phase, we design $\mathbf{X}$ as a semi-unitary matrix satisfying $\mathbf{X}^{H}\mathbf{X}=\mathbf{I}_{K}$ and set $\mathbf{w}_{k,p}$ based on the Fourier transform matrix. We model the effective antenna gain in (11) using the 3GPP standard [9].
@@
  • Fixed Parameter Scheme: In this scheme, $\{\mathbf{w}_{k}\}$, $\{\mathbf{v}_{b}\}$, $\mathbf{q}$, and $\mathbf{u}$ are fixed. We adopt a three-sector BS configuration, modeled as a special case of the IPA system with $B=3$ sectors. Each sector has approximately $\lceil\frac{NB}{3}\rceil$ antennas covering $120^{\circ}$. The precoding vector $\{\mathbf{c}_{k}\}$ uses maximum-ratio transmission (MRT), while the phase shifts and amplitudes of the polarforming vectors $\{\mathbf{w}_{k}\}$ and $\{\mathbf{v}_{b}\}$ are randomly generated and then fixed.
@@
In Fig. 11, we plot the achievable rates by different schemes versus BS’s transmit power. The results demonstrate that any level of antenna reconfiguration, i.e., proposed polarforming optimization only, position-rotation optimization only, and TT-PPT algorithm, can increase the achievable rate compared to the fixed parameter scheme, while joint antenna position/rotation and polarforming optimization by the TT-PPT algorithm achieves the highest performance gain. In particular, even the proposed polarforming optimization only scheme achieves a significant gain in achievable average rate over the fixed parameter scheme. This is due to the ability of polarforming antenna to dynamically adjust the polarization of the users and BS to maximize the instantaneous channel gain, thus effectively leveraging the additional DoFs provided by polarization diversity. Moreover, it is observed that with the same transmit power, the position-rotation optimization only scheme can also achieve higher rates than the fixed parameter scheme. This is because the IPA system with position-rotation adjustment has more spatial DoFs and can deploy the antenna resources more reasonably to match the users’ channel spatial distribution. Furthermore, the performance gap between the proposed schemes and the fixed-parameter scheme widens with increasing BS transmit power. This is expected as the sum rate becomes more interference-limited as the transmit power increases, and adjusting user/BS polarforming and the antenna positions/rotations at BS can effectively improve the multiuser-MIMO channel condition for interference suppression by BS’s transmit precoding.
@@
The effect of the number of antennas at the BS on the proposed polarforming optimization only scheme is also shown in Fig. 11. We examine two cases in which each IPA subarray at the BS employs the same polarforming vector, i.e., the proposed subarray-based polarforming scheme with $N=4$, or each antenna at the BS can have independent polarization control, i.e., the proposed antenna-based polarforming scheme with $N=1$. We observe that even though the subarray-based polarforming has less flexibility in polarforming configuration, its performance is very close to that of the antenna-based polarforming, which, however, requires higher hardware costs. This demonstrates that the proposed subarray-based polarforming scheme provides a more practically appealing balance between performance and hardware cost. Furthermore, Fig. 11 also shows that the performance advantage of polarforming becomes increasingly significant compared to the fixed parameter scheme as the number of BS antennas increases. This is because, with a larger number of BS antennas, the IPA-enhanced system possesses more polarization DoFs, which allows polarization reconfiguration to better utilize these DoFs to optimize channel conditions and improve achievable rates.
@@ VII Conclusion
In this paper, we have proposed a novel approach to enhance wireless sensing and communication performance by developing IPA, which intelligently adjusts its polarization through polarforming vector and adaptively tunes its rotation and position to cater to channel variations. Specifically, we have modeled the IPA channel and designed an efficient polarforming-based user localization algorithm that extracts the common parameters in the stable unpolarformed channel across different controllable polarforming vectors. Then, given the practical movement constraints of IPAs, the transceiver polarforming and the antenna positions/rotations were jointly optimized to maximize the weighted sum rate of users in an IPA-enhanced ISAC system. By applying the PDD and recursive sampling surrogate techniques, an efficient algorithm has been proposed to balance the system performance and computational complexity. Extensive simulation results under various practical setups demonstrated the effectiveness of localization, polarforming, and antenna placement. IPA can significantly enrich the design DoFs for sensing and communication by tuning polarforming vectors and antenna positions/rotations at a limited cost, thereby opening up promising new applications and research directions for future wireless networks.

### v2

@@ VI Simulation Results
In the simulation, we set $N=4$, $B=16$, $K=30$, and $\mathcal{C}$ as a cube with $A=1$ meter (m) sides. Users are situated within a 3D coverage area, which forms a spherical annulus ranging from $20$ m to $200$ m in radial distance from the BS center. The carrier frequency is 24 GHz. The inter-antenna spacing on each PA subarray is $\frac{\lambda}{2}$ and the minimum inter-subarray distance is $d_{\min}=\frac{\sqrt{2}}{2}\lambda+\frac{\lambda}{2}$. In Algorithm 3, we set $S=200$, $\kappa^{(i)}={i^{-0.2}}$, and $I_{\mathrm{iter}}=100$. We set the number of channel samples $\overline{L}=4000$ with $L_{\mathrm{S}}=40$, unless specified otherwise.
To start with, we evaluate the performance of the proposed polarforming-based localization algorithm and compare it with the direct positioning method in [37]. The direct positioning method sets $L=T_{\mathrm{s}}$ with $P=1$ and applies gradient‐based method to iteratively optimize the users’ DoA vectors $\mathbf{f}_{k}$, user-to-BS distances $d_{k}$, and dynamic polarformed channels $\eta_{m,k}$ to maximize the log-likelihood function. The localization error is calculated by $\sqrt{\mathbb{E}\left\{\|\hat{\mathbf{p}}-\mathbf{p}\|^{2}\right\}}$ [38], where $\mathbf{p}$ represents the true locations of all users, and $\hat{\mathbf{p}}$ denotes their estimated locations. During the localization phase, we design $\mathbf{X}$ as a semi-unitary matrix satisfying $\mathbf{X}^{H}\mathbf{X}=\mathbf{I}_{K}$ and set $\mathbf{w}_{k,p}$ based on the Fourier transform matrix. We model the effective antenna gain in (11) using the 3GPP standard [9].
@@
  • Fixed Parameter Scheme: In this scheme, $\{\mathbf{w}_{k}\}$, $\{\mathbf{v}_{b}\}$, $\mathbf{q}$, and $\mathbf{u}$ are fixed. We adopt a three-sector BS configuration, modeled as a special case of the PA system with $B=3$ sectors. Each sector has approximately $\lceil\frac{NB}{3}\rceil$ antennas covering $120^{\circ}$. The precoding vector $\{\mathbf{c}_{k}\}$ uses maximum-ratio transmission (MRT), while the phase shifts and amplitudes of the polarforming vectors $\{\mathbf{w}_{k}\}$ and $\{\mathbf{v}_{b}\}$ are randomly generated and then fixed.
@@
In Fig. 11, we plot the achievable rates by different schemes versus BS’s transmit power. The results demonstrate that any level of antenna reconfiguration, i.e., proposed polarforming optimization only, position-rotation optimization only, and TT-PPT algorithm, can increase the achievable rate compared to the fixed parameter scheme, while joint antenna position/rotation and polarforming optimization by the TT-PPT algorithm achieves the highest performance gain. In particular, even the proposed polarforming optimization only scheme achieves a significant gain in achievable average rate over the fixed parameter scheme. This is due to the ability of polarforming antenna to dynamically adjust the polarization of the users and BS to maximize the instantaneous channel gain, thus effectively leveraging the additional DoFs provided by polarization diversity. Moreover, it is observed that with the same transmit power, the position-rotation optimization only scheme can also achieve higher rates than the fixed parameter scheme. This is because the PA system with position-rotation adjustment has more spatial DoFs and can deploy the antenna resources more reasonably to match the users’ channel spatial distribution. Furthermore, the performance gap between the proposed schemes and the fixed-parameter scheme widens with increasing BS transmit power. This is expected as the sum rate becomes more interference-limited as the transmit power increases, and adjusting user/BS polarforming and the antenna positions/rotations at BS can effectively improve the multiuser-MIMO channel condition for interference suppression by BS’s transmit precoding.
@@
The effect of the number of antennas at the BS on the proposed polarforming optimization only scheme is also shown in Fig. 11. We examine two cases in which each PA subarray at the BS employs the same polarforming vector, i.e., the proposed subarray-based polarforming scheme with $N=4$, or each antenna at the BS can have independent polarization control, i.e., the proposed antenna-based polarforming scheme with $N=1$. We observe that even though the subarray-based polarforming has less flexibility in polarforming configuration, its performance is very close to that of the antenna-based polarforming, which, however, requires higher hardware costs. This demonstrates that the proposed subarray-based polarforming scheme provides a more practically appealing balance between performance and hardware cost. Furthermore, Fig. 11 also shows that the performance advantage of polarforming becomes increasingly significant compared to the fixed parameter scheme as the number of BS antennas increases. This is because, with a larger number of BS antennas, the PA-enhanced system possesses more polarization DoFs, which allows polarization reconfiguration to better utilize these DoFs to optimize channel conditions and improve achievable rates.
@@ VII Conclusion
In this paper, we have proposed a novel approach to enhance wireless sensing and communication performance by developing PA, which intelligently adjusts its polarization through polarforming vector and adaptively tunes its rotation and position to cater to channel variations. Specifically, we have modeled the PA channel and designed an efficient polarforming-based user localization algorithm that extracts the common parameters in the stable unpolarformed channel across different controllable polarforming vectors. Then, given the practical movement constraints of PAs, the transceiver polarforming and the antenna positions/rotations were jointly optimized to maximize the weighted sum rate of users in an PA-enhanced ISAC system. By applying the PDD and recursive sampling surrogate techniques, an efficient algorithm has been proposed to balance the system performance and computational complexity. Extensive simulation results under various practical setups demonstrated the effectiveness of localization, polarforming, and antenna placement. PA can significantly enrich the design DoFs for sensing and communication by tuning polarforming vectors and antenna positions/rotations at a limited cost, thereby opening up promising new applications and research directions for future wireless networks.